The Mix Cement Company sells bags of cement for $32 per bag. Fixed costs of production are $55,000 and variable costs are $10 per bag.

a. What is Mix’s break-even point in units and sales dollars?

b. How much profit will the company earn at sales levels of (i) 3,000 bags and (ii) 4,000 bags?

c. How many bags would Mix have to sell to earn a profit of $70,000?
Answer #1

(a)

Fixed Cost = $55000

Variable Cost = $10

Selling Price = $32

Let Breakeven point be x

At Breakeven point, Revenue = Cost

=> Revenue = Fixed Cost + Variable Cost

=> 32x = 55000 + 10x

=> x = 55000/(32 - 10) = 2500 units

(b) Profit = Revenue - Cost = Revenue - Fixed Cost - Variable Cost

When number of bags = 3000,
Profit = 32*3000 - 55000 - 10*3000 = $11,000

When number of bags = 4000,
Profit = 32*4000 - 55000 - 10*4000 = $33,000

(c) Let the number of bags be x

Profit = $70000

Profit = Revenue - Cost = Revenue - Fixed Cost - Variable Cost

=> 70000 = 32x - 55000 - 10x

=> x = (70000 - 55000)/(32 - 10) = 682 bags